Laramie12:

Barbara draws pens randomly from a box containing 5 pens of the same shape and size. There is 1 green pen, 3 red pens, and 1 blue pen. She draws 1 red pen and then another red pen without replacing the first one. Find the probability of drawing 1 red pen followed by another red pen, and show the equation used.

Narad:

Let total number of pens = N pens (green+ red+ blue) Let number of red pens = n red pens Probability of drawing one red pen = n/N Probability of drawing one red pen followed by another red pen (witnout replacement) \[\frac{ n }{ N } * \frac{ ( n-1) }{ (N-1) }\]

Laramie12:

what do I do now.

Narad:

You know n and N You can calculate the probability

Laramie12:

can you do that for me?

Narad:

Have a try?

Laramie12:

Ok.

Laramie12:

So we add all the number of pens and that equals= 5.

Narad:

ok and that is N

Laramie12:

Yes

Laramie12:

and for the number of red pens that is 1/5.

Narad:

How many red pens are there?

Laramie12:

one.

Narad:

No, you draw one red pen

Laramie12:

There's a total of two

Laramie12:

Since there's a another one.

Narad:

no, look at the second line

Laramie12:

its three red pens

Laramie12:

I see it now.

Narad:

what's the probability?

Laramie12:

3/5

Narad:

That's the probanility of drawing the first red pen, then how many pens are left?

Laramie12:

if you take one out.

Narad:

yes

Laramie12:

Two.

Narad:

two red pens And the total pens left are....

Laramie12:

five

Narad:

no, you have taken one (without replacement)

Laramie12:

four.

Narad:

ok, then the probability of getting the second red pen is

Laramie12:

3/4

Narad:

no

Laramie12:

2/4

Narad:

yes, Then you have the probability of drawing the first red pen and the probability of drawing the second red pen, what do you do next?

Laramie12:

Multiply

Narad:

yes

Laramie12:

3/4x2/4=6/16

Laramie12:

I think

Narad:

the first one is 3/5 and the second is 2/4

Laramie12:

3/5x2/4=6/20

Narad:

yes, you can simplify by dividing numerator and denominator by 2

Laramie12:

as a another way.

Narad:

no, simplify the final result

Laramie12:

1/3x1/2=1/6

Narad:

3/5 *2/4=3/5*1/2= ?

Laramie12:

1

Narad:

\[\frac{ 3 }{ 5 } * \frac{ 1 }{ 2 }=\]

Laramie12:

4/10

Narad:

no

Laramie12:

3/10

Narad:

yes, that's the answer

Laramie12:

Thank you.

Narad:

You are welcome